# LIGHTS. A robust technique to identify galaxy edges

### v1

@@
The LIGHTS survey is imaging galaxies at a depth and spatial resolution comparable to what the Legacy Survey of Space and Time (LSST) will produce in 10 years (i.e., 31 mag/arcsec2; 3 in areas equivalent to 10 10). This opens up the possibility of probing the edge of galaxies, as the farthest location of in-situ star formation, with a precision that we have been unable to achieve in the past. Traditionally, galaxy edges have been analyzed in one-dimension through ellipse averaging or visual inspection. Our approach allows for a two-dimensional exploration of galaxy edges, which is crucial for understanding deviations from disc symmetry and the environmental effects on galaxy growth. In this paper, we propose a novel method using the second derivative of the surface mass density map of a galaxy to determine its edges. This offers a robust quantitative alternative to traditional edge-detection methods when deep imaging is available. Our technique incorporates Wiener-Hunt deconvolution to remove the effect of the Point Spread Function (PSF) by the galaxy itself. By applying our methodology to the LIGHTS galaxy NGC 3486, we identify the edge at 205 5. At this radius, the stellar surface mass density is 1 M⊙/pc2, supporting a potential connection between galaxy edges and a threshold for in-situ star formation. Our two-dimensional analysis on NGC 3486 reveals an edge asymmetry of 5. These techniques will be of paramount importance for a physically motivated determination of the sizes of galaxies in ultra-deep surveys such as LSST, Euclid and Roman.
@@
In the framework of the Cold Dark Matter (CDM) model, galaxies form and evolve within dark matter (DM) halos. Their mass and size grow over cosmic time through both in-situ star formation, where gas within the galaxy is converted into stars, and ex-situ mechanisms, such as the accretion of stars and gas through mergers or interactions with neighboring systems (Toomre & Toomre 1972; White & Rees 1978; Efstathiou & Silk 1983). While mergers play a significant role in the growth of massive galaxies (Bundy et al. 2009), for most spirals and dwarf galaxies (see e.g., Martin et al. 2021), the primary driver of stellar mass assembly is the transformation of gas into stars, regulated by a critical gas density threshold (Quirk 1972; Kennicutt 1989; Schaye 2004; Roškar et al. 2008).
The existence of a gas density threshold should leave an imprint on a galaxy’s stellar mass radial density profile. Once the gas reservoir falls below a critical level, star formation ceases, leading to a characteristic ”drop” in the stellar mass density. The sharpness of this edge should provide insight into the recent star formation history of a galaxy. In fact, over time, stellar migration should smooth out these edges (Roškar et al. 2008; Sellwood 2014; Debattista et al. 2017), gradually erasing signatures of past star formation episodes. The prominence of a galaxy’s outer edge would eventually serve as a timer, with sharper edges indicating more recent star formation episodes (Martínez-Lombilla et al. 2018).
The critical gas density for star formation is theoretically estimated in the range of 3–10 /pc2 at z = 0 (Schaye 2004). Accounting for a star formation efficiency of 10–30, one would expect the stellar mass drop to be found at around 1–3 /pc2. This is in agreement with observational findings on present-day Milky Way-like galaxies (Martínez-Lombilla et al. 2018; Díaz-García et al. 2022).
@@
a significant advancement in understanding the gas density threshold for star formation in disc galaxies was made using a low surface brightness friendly reduction of Stripe 82 data (Trujillo & Fliri 2016; Román & Trujillo 2018), which is 2 magnitudes deeper than SDSS.
Chamba et al. (2022) revealed a moderate mass dependence, with the drop of the stellar mass profile ranging from 0.6 /pc2 in dwarf galaxies to 3 /pc2 in ellipticals.
Using the location of the 1 /pc2 isomass contour (R1; Trujillo et al. 2020) as a proxy for the star formation threshold, or the location of the stellar mass density drop as a measurement of the size of the galaxy (Redge; Chamba et al. 2022), Trujillo et al. (2020) and Chamba et al. (2022) found that the observed scatter in the stellar mass–size relation decreases by a factor larger than 2 with respect to using the effective radius, . For galaxies in the mass range , the stellar mass–size relation follows a power-law with a slope close to .
In addition, using this feature as a size measurement, Buitrago & Trujillo (2024) found, an increase of the size of the disc by a factor 2 since z for Milky Way-like disc galaxies (M).
The vast majority of photometric studies analyzing break and/or galaxy edges rely on one-dimensional (1D) surface brightness or stellar mass density profiles (Pohlen & Trujillo 2006; Bakos et al. 2008; Gutiérrez et al. 2011; Comerón et al. 2012; Martín-Navarro et al. 2012; Peters et al. 2017a, b; Trujillo et al. 2020; Chamba et al. 2022, 2024; Buitrago & Trujillo 2024). While effective, these methods average over azimuthal variations and substructures, potentially missing key morphological features. In contrast, bidimensional (2D) analyses preserve spatial information, allowing to capture asymmetries and localized angular substructures.
@@
Optimized observational strategies and data reduction methods (Jablonka et al. 2010; Trujillo & Fliri 2016; Trujillo et al. 2021; Montes et al. 2021; Román et al. 2023; Montes et al. 2024; Golini et al. 2024) now allow for surface brightness limits of mag/arcsec2 (3; 10 10 boxes) and similar depths are expected from the upcoming 10-year Legacy Survey of Space and Time (LSST; Ivezić et al. 2019). These deeper data are crucial for systematically investigate the edges of galaxies with an unprecedented accuracy.
Therefore, with the increasing volume of deep observational data, a robust, automated, and unbiased method for identifying galaxy edges is essential, as previous studies have primarily relied on visual detection.
In computer vision, boundaries are identified by detecting rapid transitions in brightness, color, or texture (see Jing et al. 2022 for a review). However, defining a galaxy’s edge requires a physically motivated approach rather than purely visual cues. While some methods, such as fitting intersecting exponential functions (Martínez-Lombilla et al. 2023), have been proposed to objectively locate edges, a fully automated approach is crucial for ensuring consistency across large datasets.
A recent study by Fernández-Iglesias et al. (2024) applied artificial intelligence (AI) to automate the identification of galaxy edges using Hubble images. By analyzing different combinations of color images, their algorithm successfully recovered previously visually identified edges (Buitrago & Trujillo 2024). Similarly, Vega-Ferrero et al. (2025) apply the Segment Anything Model (SAM) to recover the radial positions of the stellar disk edges from Buitrago & Trujillo (2024).
@@
The goal of our work is to present a robust method for identifying galaxy edges by locating the minimum of the second derivative of the stellar mass density distribution. This technique provides a reproducible approach to measuring galaxy sizes across large datasets, facilitating a more systematic study of their outermost regions. Additionally, it will enable the investigation of the stellar mass density threshold at the edge across different morphologies, environments (Chamba et al. 2024), and will allow for the exploration of the cosmic evolution of galaxy sizes (Buitrago & Trujillo 2024) with the upcoming next generation datasets.
This paper is structured as follows: Section 2 describes the methodology to locate the edge of galaxies. Section 3 presents the LBT Imaging of Galaxy Haloes and Tidal Structures (LIGHTS; Trujillo et al. 2021; Zaritsky et al. 2024) data used in this work for the galaxy NGC 3486. Section 4 details data processing and mitigation of scattered light effects. Sections 5 and 6 discuss our findings and summarize conclusions. All magnitudes are reported in the AB system (Oke & Gunn 1983).
@@ 2 A robust procedure to locate galaxy edges
We propose a method that uses the second derivative of the stellar mass density distribution, , to determine the galaxy edge. By finding the radial position at which the second derivative, reaches the outermost111The term ”outermost” is used because inner regions of a galaxy, such as the bulge, bar transition, spiral arms, or clumps, can also produce additional minima. minimum, we identify the point of curvature corresponding to the outer sharp change in slope in . This point marks a natural transition in the galaxy’s structure and provides a robust, identification of its edge.
@@ 2.1 Motivation for using the stellar mass density profile
If a gas density threshold is indeed present (Schaye 2004; Roškar et al. 2008) affecting star formation, both surface brightness and stellar mass density profiles should reflect a structural transition: one in terms of light () and the other in terms of mass (). In this work we use the stellar mass density profile. The reason lies in the nature of the stellar populations. In surface brightness profiles, particularly in bluer bands, variations in the luminosity of young star-forming regions, spiral arms, and clumps introduce localized fluctuations, which can create multiple substructures in the profile. These can blur the detection of the outermost drop in the averaged lights profile, as they introduce local minima of the second derivative. In contrast, the stellar mass density profile, derived from the full Spectral Energy Distribution (SED) or from color combinations, smooths out variations in stellar populations across the disc, reducing the influence of substructures that are more prominent in light-based profiles (see Fig. 8). Working with profiles has also some drawbacks. The most important one is the addition of larger uncertainties as it relies on the combination of multiple surface brightness profiles plus the adoption of some stellar population models. For these reasons, the ideal scenario would be to work with ultra-deep Near-Infrared (NIR) data, where the underlying mass distributions can be traced more directly. In this work we rely on Sloan and Sloan data to get the stellar mass density profile of galaxies, as these are the available filters in the LIGHTS survey.
@@ 2.2 Mathematical derivation
As mentioned in the Introduction, the end of in-situ star formation should lead to a drop in the stellar mass distribution of a galaxy disc. Therefore, in the majority of galaxy discs and dwarfs, the mass distribution should be well represented by a broken exponential function (a function described in Erwin et al. 2008) that has two defined exponential declining zones with different scale lengths. We define the transition point as the edge of the galaxy .
Within the galaxy edge (r ¡ ), the slope of the disc’s stellar mass surface density profile, , is parametrized by the scale length . Beyond , the profile declines with a second scale length . The sharpness of the edge is parametrized by . For an infinitely thin disc, the broken exponential is described as:
The first derivative with respect to the radial position indicates how the intensity changes along the radius. For a broken exponential function in logarithmic scale , the first derivative shows a sharp change at the point where the slope changes, , preceded and followed by two constants (equal to - 1/ before and to - 1/ beyond). The second derivative gives information about the curvature or the rate of change of the slope. At the point where the slope changes abruptly, the second derivative has a minimum. Therefore, presents a minimum corresponding to . In other words, the search for the minimum of corresponds to the search for the location of .
2.3 Application to a simulated disc
To illustrate the implementation of our second-derivative method for edge detection, we generated a simulated disc using IMFIT (Erwin 2015), an image-fitting tool designed for modeling galaxy surface brightness profiles. The disc follows a BrokenExponential light distribution, with an inner scale length of = 45 pixels transitioning to a steeper declining outer slope of = 11 pixels at = 200 pixels. We assume a sharp edge with = 10, Position Angle (PA) = 30 deg (counterclockwise from the X axis) and axis ratio (AR) of 0.7.
The top-left panel of Fig. 1 shows the stellar mass density distribution of the simulated disc in Cartesian coordinates, where the black dashed ellipse marks the point of change of slope (). The top-central panel presents the same distribution in Polar coordinates and, the sinusoidal wave corresponding to the edge location which allows to examine possible asymmetries in the edge structure.
@@ 2.3 Application to a simulated disc
The bottom row of Fig. 1 summarizes the one-dimensional analysis, showing (from left to right) the stellar mass density profile, its first derivative, and second derivative in three different cases: the analytical BrokenExponential function used to generate the simulated disc with IMFIT, the averaged stellar mass surface density profile of the disc, and the 30-degree222A high-degree polynomial capture subtle variations and provides high fitting precision. polynomial that fits the averaged stellar mass surface density profile of the disc. The averaged stellar mass surface density profile of the disc is extracted using elliptical apertures with fixed position angle and inclination, and a radial step size of one pixel with astscript-radial-profile in Gnuastro (Infante-Sainz et al. 2024).
The derivatives of the BrokenExponential function and the polynomial fit are calculated analytically. The derivatives of the averaged stellar mass density profile of the simulated disc are computed using a custom implementation of the five-point-stencil derivative333The error in this approximation of derivatives is of order 4 (Abramowitz & Stegun 1948). method in Python (Sauer 2011). This numerical method leverages a higher-order Taylor series expansion, yielding a more precise derivative by drawing on multiple neighboring points. The approach reduces local errors and smooths out irregularities, though it does cause the loss of the first and last data points. However, this is inconsequential here, as our primary focus is on the outer regions of the galaxy, well-covered within our data range. The edge location remains consistent in the three cases, demonstrating the robustness of our method using simulated data. The minimum of the second derivative corresponds to (dashed red lines in Fig. 1).
@@
While fixed elliptical apertures are a standard method for surface photometry, they average over the entire azimuthal range, potentially smoothing out localized variations. In contrast, wedge444Wedges are a sector of a circle that is defined by two straight lines (radii) extending from the center of the outer regions, and the arc connecting them.-based profiles allow for a more localized analysis, capturing directional differences and small-scale features, even though they still require assumptions about the geometry in each sector.
The elliptical symmetry of the simulated disc ensures that the minimum of the second derivative along each radial direction, , corresponds to by design. The top-right panel of Fig. 1 presents the second derivative map in Polar coordinates555In this image, each column is treated as a one-dimensional array, and the pixel values represent the numerical second derivatives computed along each column., where a clear sinusoidal pattern emerges (in pink), corresponding to the edge. This structure marks the edge across different azimuthal directions. The dispersion between gray lines in the bottom panels of Fig. 1 represents the uncertainty in computing the derivatives using 10 degrees wedges along different azimuthal angles. No error bars are shown, as we assume an idealized scenario with infinite signal to noise (S/N). Despite the disc model being noise-free, minor inaccuracies in locating R and artificial oscillations of the derivatives, arise from numerical artifacts introduced during the discrete generation of the 2D image, and radial or angular binning of the profile. The choice of the width of the wedge, depending on the S/N of real data, is discussed in next sections.

### v2

@@
The LIGHTS survey is imaging galaxies at a depth and spatial resolution comparable to what the Legacy Survey of Space and Time (LSST) will produce in 10 years (i.e., 31 mag/arcsec2; 3 in areas equivalent to 10 10). This opens up the possibility of probing the edges of galaxies, as the farthest location of in situ star formation, with a precision that we have been unable to achieve in the past. Traditionally, galaxy edges have been analyzed in one dimension through ellipse averaging or visual inspection. Our approach allows for a two-dimensional exploration of galaxy edges, which is crucial for understanding deviations from disk symmetry and the environmental effects on galaxy growth. In this paper, we propose a novel method using the second derivative of the surface mass density map of a galaxy to determine its edges. This offers a robust quantitative alternative to traditional edge-detection methods when deep imaging is available. Our technique incorporates Wiener-Hunt deconvolution to remove the effect of the point spread function from the galaxy itself. By applying our methodology to the LIGHTS galaxy NGC 3486, we identify the edge at 205 5. At this radius, the stellar surface mass density is 1 M⊙/pc2, supporting a potential connection between galaxy edges and a threshold for in situ star formation. Our two-dimensional analysis of NGC 3486 reveals an edge asymmetry of 5. These techniques will be of paramount importance for a physically motivated determination of the sizes of galaxies in ultra-deep surveys such as LSST, Euclid, and Roman.
@@
In the framework of the cold dark matter (CDM) model, galaxies form and evolve within dark matter (DM) halos. Their mass and size grow over cosmic time through both in situ star formation, whereby gas within the galaxy is converted into stars, and ex situ mechanisms, such as the accretion of stars and gas through mergers or interactions with neighboring systems (Toomre & Toomre 1972; White & Rees 1978; Efstathiou & Silk 1983). While mergers play a significant role in the growth of massive galaxies (Bundy et al. 2009), for most spirals and dwarf galaxies (see e.g., Martin et al. 2021) the primary driver of stellar mass assembly is the transformation of gas into stars, regulated by a critical gas density threshold (Quirk 1972; Kennicutt 1989; Schaye 2004; Roškar et al. 2008).
The existence of a gas density threshold should leave an imprint on a galaxy’s stellar mass radial density profile. Once the gas reservoir falls below a critical level, star formation ceases, leading to a characteristic “drop” in the stellar mass density. The sharpness of this edge should provide an insight into the recent star formation history of a galaxy. In fact, over time, stellar migration should smooth out these edges (Roškar et al. 2008; Sellwood 2014; Debattista et al. 2017), gradually erasing signatures of past star formation episodes. The prominence of a galaxy’s outer edge would eventually serve as a timer, with sharper edges indicating more recent star formation episodes (Martínez-Lombilla et al. 2018).
The critical gas density for star formation is theoretically estimated in the range of 3–10 /pc2 at z = 0 (Schaye 2004). Accounting for a star formation efficiency of 10–30, one would expect the stellar mass drop to be found at around 1–3 /pc2. This is in agreement with observational findings for present-day Milky Way-like galaxies (Martínez-Lombilla et al. 2018; D´ıaz-Garc´ıa et al. 2022).
@@
a significant advancement in understanding the gas density threshold for star formation in disk galaxies was made using a low-surface-brightness (LSB) friendly reduction of Stripe 82 data (Trujillo & Fliri 2016; Román & Trujillo 2018), which is 2 magnitudes deeper than SDSS.
Chamba et al. (2022) revealed a moderate mass dependence, with the drop in the stellar mass profile ranging from 0.6 /pc2 in dwarf galaxies to 3 /pc2 in ellipticals.
Using the location of the 1 /pc2 isomass contour (R1; Trujillo et al. 2020) as a proxy for the star formation threshold, or the location of the stellar mass density drop as a measurement of the size of the galaxy (Redge; Chamba et al. 2022), Trujillo et al. (2020) and Chamba et al. (2022) found that the observed scatter in the stellar mass–size relation decreases by a factor larger than 2 with respect to using the effective radius, . For galaxies in the mass range , the stellar mass–size relation follows a power law with a slope close to .
In addition, using this feature as a size measurement, Buitrago & Trujillo (2024) found an increase in the size of the disk by a factor of 2, since z for Milky Way-like disk galaxies (M).
The vast majority of photometric studies analyzing breaks and/or galaxy edges rely on one-dimensional (1D) surface brightness or stellar mass density profiles (Pohlen & Trujillo 2006; Bakos et al. 2008; Gutiérrez et al. 2011; Comerón et al. 2012; Mart´ın-Navarro et al. 2012; Peters et al. 2017a, b; Trujillo et al. 2020; Chamba et al. 2022, 2024; Buitrago & Trujillo 2024). While effective, these methods average over azimuthal variations and substructures, potentially missing key morphological features. In contrast, bidimensional (2D) analyses preserve spatial information, allowing to capture asymmetries and localized angular substructures.
@@
Optimized observational strategies and data reduction methods (Jablonka et al. 2010; Trujillo & Fliri 2016; Trujillo et al. 2021; Montes et al. 2021; Román et al. 2023; Montes et al. 2024; Golini et al. 2024) now allow for surface brightness limits of mag/arcsec2 (3; 10 10 boxes) and similar depths are expected from the upcoming 10-year Legacy Survey of Space and Time (LSST; Ivezić et al. 2019). These deeper data are crucial for systematically investigating the edges of galaxies with an unprecedented accuracy.
Therefore, with the increasing volume of deep observational data, a robust, automated, and unbiased method of identifying galaxy edges is essential, as previous studies have primarily relied on visual detection.
In computer vision, boundaries are identified by detecting rapid transitions in brightness, color, or texture (see Jing et al. 2022 for a review). However, defining a galaxy’s edge requires a physically motivated approach rather than purely visual cues. While some methods, such as fitting intersecting exponential functions (Mart´ınez-Lombilla et al. 2023), have been proposed to objectively locate edges, a fully automated approach is crucial for ensuring consistency across large datasets.
A recent study by Fernández-Iglesias et al. (2024) applied artificial intelligence to automate the identification of galaxy edges using Hubble images. By analyzing different combinations of color images, their algorithm successfully recovered previously visually identified edges (Buitrago & Trujillo 2024). Similarly, Vega-Ferrero et al. (2025) applied the segment anything model to recover the radial positions of the stellar disk edges from Buitrago & Trujillo (2024).
@@
The goal of our work is to present a robust method of identifying galaxy edges by locating the minimum of the second derivative of the stellar mass density distribution. This technique provides a reproducible approach to measuring galaxy sizes across large datasets, facilitating a more systematic study of their outermost regions. Additionally, it will enable the investigation of the stellar mass density threshold at the edge across different morphologies, environments (Chamba et al. 2024), and will allow for the exploration of the cosmic evolution of galaxy sizes (Buitrago & Trujillo 2024) with the upcoming next generation of datasets.
This paper is structured as follows. Section 2 describes the methodology to locate the edge of galaxies. Section 3 presents the LBT Imaging of Galaxy Haloes and Tidal Structures (LIGHTS; Trujillo et al. 2021; Zaritsky et al. 2024) data used in this work for the galaxy NGC 3486. Section 4 details data processing and mitigation of scattered light effects. Sections 5 and 6 discuss our findings and summarize conclusions. All magnitudes are reported in the AB system (Oke & Gunn 1983).
@@ 2 A robust procedure to locate galaxy edges
We propose a method that uses the second derivative of the stellar mass density distribution, , to determine a galaxy’s edge. By finding the radial position at which the second derivative reaches the outermost111The term “outermost” is used because inner regions of a galaxy, such as the bulge, bar transition, spiral arms, or clumps, can also produce additional minima. minimum, we identify the point of curvature corresponding to the outer sharp change in slope in . This point marks a natural transition in the galaxy’s structure and provides a robust identification of its edge.
@@ 2.1 Motivation for using the stellar mass density profile
If a gas density threshold is indeed present (Schaye 2004; Roškar et al. 2008) affecting star formation, both surface brightness and stellar mass density profiles should reflect a structural transition: one in terms of light () and the other in terms of mass (). In this work we use the stellar mass density profile. The reason lies in the nature of the stellar populations. In surface brightness profiles, particularly in bluer bands, variations in the luminosity of young star-forming regions, spiral arms, and clumps introduce localized fluctuations that can create multiple substructures in the profile. These can blur the detection of the outermost drop in the averaged lights profile, as they introduce local minima of the second derivative. In contrast, the stellar mass density profile, derived from the full spectral energy distribution or from color combinations, smooths out variations in stellar populations across the disk, reducing the influence of substructures that are more prominent in light-based profiles (see Fig. 8). Working with profiles also has some drawbacks. The most important one is the addition of larger uncertainties as it relies on the combination of multiple surface brightness profiles plus the adoption of some stellar population models. For these reasons, the ideal scenario would be to work with ultra-deep near-infrared data, for which the underlying mass distributions can be traced more directly. In this work we rely on Sloan and Sloan data to get the stellar mass density profile of galaxies, as these are the available filters in the LIGHTS survey.
@@ 2.2 Mathematical derivation
As was mentioned in the introduction, the end of in situ star formation should lead to a drop in the stellar mass distribution of a galaxy disk. Therefore, in the majority of galaxy disks and dwarfs, the mass distribution should be well represented by a broken exponential function (a function described in Erwin et al. 2008) that has two defined exponential declining zones with different scale lengths. We define the transition point as the edge of the galaxy, .
Within the galaxy edge (r ¡ ), the slope of the disk’s stellar mass surface density profile, , is parametrized by the scale length, . Beyond , the profile declines with a second scale length, . The sharpness of the edge is parametrized by . For an infinitely thin disk, the broken exponential is described as
The first derivative, , with respect to the radial position indicates how the intensity changes along the radius. For a broken exponential function on a logarithmic scale, , the first derivative, , shows a sharp change at the point where the slope changes, , preceded and followed by two constants (equal to - 1/ before and to - 1/ beyond). The second derivative, , gives information about the curvature or the rate of change in the slope. At the point where the slope changes abruptly, the second derivative has a minimum. Therefore, presents a minimum corresponding to . In other words, the search for the minimum of corresponds to the search for the location of .
2.3 Application to a simulated disk
To illustrate the implementation of our second-derivative method of edge detection, we generated a simulated disk using IMFIT (Erwin 2015), an image-fitting tool designed for modeling galaxy surface brightness profiles. The disk follows a BrokenExponential light distribution, with an inner scale length of = 45 pixels transitioning to a steeper declining outer slope of = 11 pixels at = 200 pixels. We assume a sharp edge with = 10, a position angle (PA) of 30 deg (counterclockwise from the X axis), and an axis ratio (AR) of 0.7.
The top left panel of Fig. 1 shows the stellar mass density distribution of the simulated disk in Cartesian coordinates, where the dashed ellipse marks the point of change in the slope (). The top central panel presents the same distribution in polar coordinates and the sinusoidal wave corresponding to the edge location, which allows one to examine possible asymmetries in the edge structure.
@@ 2.3 Application to a simulated disk
The bottom row of Fig. 1 summarizes the 1D analysis, showing (from left to right) the stellar mass density profile, its first derivative, and its second derivative in three different cases: the analytical BrokenExponential function used to generate the simulated disk with IMFIT, the averaged stellar mass surface density profile of the disk, and the 30-degree222A high-degree polynomial capture subtle variations and provides high fitting precision. polynomial that fits the averaged stellar mass surface density profile of the disk. The averaged stellar mass surface density profile of the disk was extracted using elliptical apertures with a fixed PA and inclination, and a radial step size of one pixel with the astscript-radial-profile in Gnuastro (Infante-Sainz et al. 2024).
The derivatives of the BrokenExponential function and the polynomial fit were calculated analytically. The derivatives of the averaged stellar mass density profile of the simulated disk were computed using a custom implementation of the five-point-stencil derivative333The error in this approximation of derivatives is of order 4 (Abramowitz & Stegun 1948). method in Python (Sauer 2011). This numerical method leverages a higher-order Taylor series expansion, yielding a more precise derivative by drawing on multiple neighboring points. The approach reduces local errors and smooths out irregularities, though it does cause the loss of the first and last data points. However, this is inconsequential here, as our primary focus is on the outer regions of the galaxy, which are well covered within our data range. The edge location remains consistent in the three cases, demonstrating the robustness of our method using simulated data. The minimum of the second derivative corresponds to (dashed red lines in Fig. 1).
@@
While fixed elliptical apertures are a standard method in surface photometry, they average over the entire azimuthal range, potentially smoothing out localized variations. In contrast, wedge444A wedge is a sector of a circle that is defined by two straight lines (radii) extending from the center of the outer regions, and the arc connecting them.-based profiles allow for a more localized analysis, capturing directional differences and small-scale features, even though they still require assumptions about the geometry in each sector.
The elliptical symmetry of the simulated disk ensures that the minimum of the second derivative along each radial direction, , corresponds to by design. The top right panel of Fig. 1 presents the second derivative map in polar coordinates555In this image, each column is treated as a 1D array, and the pixel values represent the numerical second derivatives computed along each column., where a clear sinusoidal pattern emerges (in pink), corresponding to the edge. This structure marks the edge in different azimuthal directions. The dispersion between gray lines in the bottom panels of Fig. 1 represents the uncertainty in computing the derivatives using 10 degrees wedges along different azimuthal angles. No error bars are shown, as we assume an idealized scenario with infinite signal to noise (S/N). Despite the disk model being noise-free, minor inaccuracies in locating R and artificial oscillations of the derivatives arise from numerical artifacts introduced during the discrete generation of the 2D image, and radial or angular binning of the profile. The choice of width of the wedge, depending on the S/N of real data, is discussed in the next sections.